2v + 18 = 16 - 4(v + 7)
Answer by checkley77(12844)

2v + 18 = 16 - 4(v + 7)
2v+18=16-4v-28
2v+4v=16-28-18
6v=-30
v=-30/6
v=-5 answer.
Proof:
2(-5)+18=16-4(-5+7)
-10+18=16-4*2
8=16-8
8=8